OpenStudy (anonymous):

(xto the 9th power y to the 5th power)all to -4 power

OpenStudy (gorv):

x^9*y^5)^-4

OpenStudy (anonymous):

yes that's the question? How do i solve it

OpenStudy (gorv):

x^(9*(-4))*y^5*(-4)=x^(-36)*y^(-20)

OpenStudy (gorv):

just take the power -4 inside and multiply with the powers thatzz it

OpenStudy (anonymous):

i have to get rid of negative exponets

OpenStudy (anonymous):

would you put it all under 1

OpenStudy (gorv):

okk \[a^{-b}=\frac{ 1 }{ a^b }\]

OpenStudy (anonymous):

1 over x^36y^20

OpenStudy (gorv):

yeah you r right

OpenStudy (gorv):

itzz correct

OpenStudy (anonymous):

thank you , i have more to come lol

OpenStudy (anonymous):

(negative 1 over 4 ) to -4